# RetroInfer: A Vector-Storage Approach for Scalable Long-Context LLM Inference

### v1

@@ 5.1. Experimental Setup
We run experiments on a VM server with an NVIDIA A100 GPU (80GB memory) and an AMD EPYC 7V12 64-core CPU (1900GB memory). The VM has 4 NUMA nodes, each with 12 physical cores. The GPU and CPU are linked via PCIe 4.0 (×$\times 16$), with a unidirectional bandwidth of 32GB/s. The server runs Ubuntu 22.04 with CUDA 12.4 and PyTorch 2.4.0.
Models. We evaluate RetroInfer on three leading open-source LLMs: Llama 3.1-8B (llama3.1, ), Qwen2.5-7B (qwen2.5, ), and Llama3-8B-1048K (llama3-8B-1048k, ). Llama 3.1-8B (32 layers, 32 heads) and Qwen2.5-7B (28 layers, 28 heads) support up to 128K tokens. These models allow evaluation across architectures. All models adopt GQA, with group size 4 for Llama and 7 for Qwen. Llama3-8B-1048K, architecturally identical to Llama3.1-8B but with a 1048K context window, demonstrates effectiveness under extreme long-context scenarios.
@@
Parameters. Each sparse attention baseline uses the hyperparameters from its original paper  [^4] Quest uses a chunk size of 16; MagicPIG employs 10 hash functions and 150 hash tables per KV head; InfiniGen uses 32 partial channels. and RetroInfer uses $\frac{1}{16}$ centroids of all tokens (avg. cluster size 16), with segments of 8K tokens and 10 iterations for $k$-means. For Quest, full attention is applied to the first two layers. For MagicPIG, full attention is applied to layer-0, 16 of Llama3-8B and Llama3.1-8B, and layer-0, 14 of Qwen2.5-7B.
@@ 5.2. Inference Accuracy
Table 1. Model Accuracy (RULER, 128K Context). RetroInfer significantly outperforms other methods and maintains the same accuracy level as full attention on three models.
@@
Table 2. Model Accuracy (LongBench). RetroInfer consistently outperforms all baselines and maintains the same accuracy level as full attention on three models.
Methods QaS GoR TrQ RbP LCC Avg.
Llama3.1-8B 12.88 34.30 91.31 56.40 62.88 51.55
Quest 9.43 31.42 89.86 50.41 44.30 45.08
MagicPIG 12.00 33.18 90.80 55.50 61.14 50.52
InfiniGen 10.96 32.36 90.71 49.94 44.12 45.62
RetroInfer 12.31 34.71 91.33 56.15 62.10 51.32
Qwen2.5-7B 9.98 35.41 87.15 61.69 56.34 50.11
Quest 9.05 31.54 82.16 53.15 42.07 43.59
MagicPIG 9.87 34.00 86.35 60.25 55.43 49.18
InfiniGen 9.61 33.10 87.02 46.72 30.58 41.41
RetroInfer 9.77 34.76 87.33 61.80 55.76 49.88
Llama3-8B-1048K 14.25 35.25 87.06 44.74 45.18 45.30
Quest 11.22 30.35 86.94 44.84 36.74 42.02
MagicPIG 13.66 33.28 87.21 42.56 43.58 44.06
InfiniGen 12.15 30.81 85.39 38.84 29.41 39.32
RetroInfer 15.17 34.66 86.74 45.58 45.66 45.56
Different Tasks. Table 1 shows the model accuracy of various methods on the RULER benchmark with 128K context length. RetroInfer consistently achieves better accuracy than other methods on almost all tasks and matches the accuracy of full attention. For the average task accuracy, RetroInfer has only a drop of 0.73%/0.78%/1.46% compared to full attention on Llama3.1-8B/Qwen2.5-7B/Llama3-8B-1048K, respectively. Compared to the best-performing sparse attention baselines on each model (Quest/InfiniGen/Quest for Llama3.1-8B/Qwen2.5-7B/Llama3-8B-1048K), RetroInfer achieves 5.48%/15.51%/3.33% accuracy improvements, respectively. The accuracy improvements over other methods are even larger, up to 17.02%/19.51%/6.26% on three models. These results indicate that RetroInfer is able to effectively retrieve important tokens using wave index and cover the varying sparsity by attention estimation. In certain tasks (e.g., fwe on model Llama3.1-8B), RetroInfer even outperforms full attention, possibly because it helps models focus on important tokens and filter out distractions.
To further evaluate the effectiveness of RetroInfer on more realistic tasks, we selected five tasks from the LongBench benchmark for extensive evaluation. It includes single-document QA (Qasper), summarization (GovReport), few-shot learning (TriviaQA), code completion (RepoBench-P) and LCC. As shown in Table 2, RetroInfer consistently outperforms all baselines, with average scores at most 0.23% lower than full attention.
@@ 5.3. Inference Efficiency
We also evaluate the decoding throughput on different tasks and models. As shown in Figure 13 (a), RetroInfer outperforms full attention by 4.4$\times$, 4.0$\times$, 4.6$\times$, and 3.4$\times$ on NIAH, fwe, qa1 and vt tasks, respectively. The throughput variation across tasks is due to differing cache hit ratios, as we analyzed later ( §5.4). The advantage over sparse attention systems still holds, with RetroInfer achieving 3.3$\times$, 3.0$\times$, 3.5$\times$, and 2.6$\times$ higher throughput than the best-performing baseline Quest on four tasks. Figure 13 (b) shows the decoding throughput of different models. RetroInfer outperforms full attention by 3.6$\times$, 2.9$\times$, and 4.4$\times$ on Llama3.1-8B, Qwen2.5-7B and Llama3-8B-1048K, respectively. Although Llama3.1-8B has the same model architecture as Llama3-8B-1048K, it has throughput difference because of different cache behaviors. For Qwen2.5-7B, with less memory consumption due to fewer KV heads and model layers, the advantages of RetroInfer are still evident, demonstrating its architectural generality.

### v2

@@ 4.2. Wave Index
Lightweight Index Update. During the decoding phase, following segmented clustering, the index is updated incrementally whenever the generated tokens exceeds a segment. Rather than rebuilding the entire index, we perform $k$-means clustering only on the newly generated tokens, significantly reducing the index update overhead. This design leverages the coarse-grained spatial locality observed in the generated key vectors, consistent with patterns seen during the prefilling stage. The clustering results from the new segment are appended to the existing index, while the corresponding KV vectors are offloaded to the CPU. Experimental results in §5.5 demonstrate the incremental index update is both efficient and capable of preserving model accuracy.
@@ 5.1. Experimental Setup
We run experiments on a VM server with NVIDIA A100 GPUs (80GB memory) and an AMD EPYC 7V12 64-core CPU (1900GB memory). The VM has 4 NUMA nodes, each with 12 physical cores. The GPU and CPU are linked via PCIe 4.0 (×$\times 16$), with a unidirectional bandwidth of 32GB/s. The server runs Ubuntu 22.04 with CUDA 12.4 and PyTorch 2.4.0.
Models. We evaluate RetroInfer on four leading open-source LLMs: Llama 3.1-8B (llama3.1, ), Qwen2.5-7B (qwen2.5, ), Llama3-8B-1048K (llama3-8B-1048k, ) and Qwen2.5-72B (qwen2.5-72b, ). Llama 3.1-8B (32 layers, 32 heads), Qwen2.5-7B (28 layers, 28 heads)and Qwen2.5-72B (80 layers, 64 heads) support up to 128K tokens. These models allow evaluation across architectures. All models adopt GQA, with group size 4 for Llama, 7 for Qwen-7B and 8 for Qwen-72B. Llama3-8B-1048K, architecturally identical to Llama3.1-8B but with a 1048K context window, demonstrates effectiveness under extreme long-context scenarios. Unless otherwise specified, all models are evaluated on a single A100 GPU except that Qwen2.5-72B is evaluated on eight A100 GPUs by partitioning the model layers across GPUs due to its large memory footprint. The evaluation on Qwen2.5-72B stress tests RetroInfer’s scalability to large models and multiple GPUs. Section 5.5 evaluates RetroInfer’s performance on long-context generation tasks using reasoning models.
@@
Parameters. Each sparse attention baseline uses the hyperparameters from its original paper  [^4] Quest uses a chunk size of 16; MagicPIG employs 10 hash functions and 150 hash tables per KV head; InfiniGen uses 32 partial channels. and RetroInfer uses $\frac{1}{16}$ centroids of all tokens (avg. cluster size 16), with segments of 8K tokens and 10 iterations for $k$-means. For Quest, full attention is applied to the first two layers. For MagicPIG, full attention is applied to layer-0, 16 of Llama3-8B and Llama3.1-8B, layer-0, 14 of Qwen2.5-7B, and layer-0, 20, 40, 60 for Qwen2.5-72B.
@@ 5.2. Inference Accuracy
Table 1. Model Accuracy (RULER, 128K Context). RetroInfer significantly outperforms other methods and maintains the same accuracy level as full attention on four models.
@@
Qwen2.5-72B 100.00 99.00 98.50 97.00 86.00 37.50 96.50 98.50 87.17 89.60 53.00 48.50 99.70 83.92
Quest 100.00 92.00 92.50 66.50 30.50 1.50 68.12 80.75 84.67 45.45 29.00 30.50 84.80 62.02
MagicPIG 100.00 94.50 96.50 95.00 85.50 27.00 92.12 92.12 86.50 89.75 49.50 47.50 99.60 81.20
InfiniGen 100.00 99.00 100.00 95.50 80.50 21.00 96.50 98.50 86.33 71.15 54.50 49.00 99.60 80.89
RetroInfer 100.00 99.50 100.00 97.00 87.50 32.00 94.88 98.25 84.83 93.35 52.50 48.00 99.60 83.65
Table 2. Model Accuracy (LongBench). RetroInfer consistently maintains almost the same accuracy level as full attention on four models.
Methods SQA MQA SUM FL ST CC Avg.
Llama3.1-8B 13.81 17.82 25.24 63.52 60.35 59.64 38.40
Quest 12.00 15.61 23.22 61.33 51.92 46.57 34.27
MagicPIG 13.36 16.98 24.67 62.73 60.18 58.48 37.75
InfiniGen 12.65 15.86 22.97 61.82 54.14 46.95 34.84
RetroInfer 13.83 17.94 24.84 62.96 61.23 59.18 38.32
Qwen2.5-7B 12.79 15.48 23.69 60.57 56.39 59.10 36.23
Quest 11.75 13.37 22.00 57.27 49.46 47.06 32.41
MagicPIG 12.74 14.08 23.44 57.99 43.92 57.18 33.32
InfiniGen 12.53 14.58 21.35 54.94 41.02 38.59 30.06
RetroInfer 12.62 15.28 23.18 60.13 56.06 57.24 35.74
Llama3-8B-1048K 17.62 14.84 23.41 59.76 42.37 44.98 33.10
Quest 14.09 15.36 20.10 58.09 36.43 39.87 30.27
MagicPIG 16.05 14.28 22.60 57.99 38.10 43.81 31.48
InfiniGen 15.85 16.05 21.25 56.37 31.40 34.07 29.23
RetroInfer 17.39 15.81 22.41 59.51 41.08 45.56 32.87
Qwen2.5-72B 13.23 18.46 23.59 66.41 36.15 51.28 39.70
Quest 13.18 17.08 22.42 65.66 35.05 50.35 38.06
MagicPIG 13.39 18.27 23.38 66.51 36.05 51.28 39.36
InfiniGen 13.48 17.98 23.91 66.18 36.02 51.10 39.00
RetroInfer 13.26 18.20 23.38 66.20 35.93 51.06 39.15
Different Tasks. Table 1 shows the model accuracy of various methods on the RULER benchmark with 128K context length. RetroInfer consistently achieves better accuracy than other methods on almost all tasks and matches the accuracy of full attention. For the average task accuracy, RetroInfer has only a drop of 0.73%/0.78%/1.46%/0.27% compared to full attention on Llama3.1-8B/Qwen2.5-7B/Llama3-8B-1048K/Qwen2.5-72B, respectively. Compared to the best-performing sparse attention baselines on each model (Quest/InfiniGen/Quest/MaigcPIG for Llama3.1-8B/Qwen2.5-7B/Llama3-8B-1048K/Qwen2.5-72B), RetroInfer achieves 5.48%/15.51%/3.33%/2.45% accuracy improvements, respectively. The accuracy improvements over other methods are even larger, up to 17.02%/19.51%/6.26%/21.63% on four models. These results indicate that RetroInfer is able to effectively retrieve important tokens using wave index and cover the varying sparsity by attention estimation. In certain tasks (e.g., fwe on model Llama3.1-8B), RetroInfer even outperforms full attention, possibly because it helps models focus on important tokens and filter out distractions.
To further evaluate the effectiveness of RetroInfer on more realistic tasks, we selected the LongBench benchmark for extensive evaluation. It includes single-document QA (SQA), multi-document QA (MQA), summarization (SUM), few-shot learning (FL), synthetic task(ST) and code completion (CC). As shown in Table 2, RetroInfer consistently outperforms all baselines, with average scores at most 0.23% lower than full attention.
@@ 5.3. Inference Efficiency
We also evaluate the decoding throughput on different tasks and models. As shown in Figure 13 (a), RetroInfer outperforms full attention by 4.4$\times$, 4.0$\times$, 4.6$\times$, and 3.4$\times$ on NIAH, fwe, qa1 and vt tasks, respectively. The throughput variation across tasks is due to differing cache hit ratios, as we analyzed later ( §5.4). The advantage over sparse attention systems still holds, with RetroInfer achieving 3.3$\times$, 3.0$\times$, 3.5$\times$, and 2.6$\times$ higher throughput than the best-performing baseline Quest on four tasks. Figure 13 (b) shows the decoding throughput of different models. RetroInfer outperforms full attention by 3.6$\times$, 2.9$\times$, 4.4$\times$ and 2.8$\times$ on Llama3.1-8B, Qwen2.5-7B, Llama3-8B-1048K and Qwen2.5-72B, respectively. Although Llama3.1-8B has the same model architecture as Llama3-8B-1048K, it has throughput difference because of different cache behaviors. For Qwen2.5-7B, with less memory consumption due to fewer KV heads and model layers, the advantages of RetroInfer are still evident, demonstrating its architectural generality. The results on Qwen2.5-72B demonstrate that RetroInfer can scale effectively to multi-GPU setups and larger models. Note that the performance of RetroInfer on Qwen2.5-72B is constrained by the CPU memory capacity of the test server; with more available CPU memory, even greater throughput gains would be achieved.
@@ 5.4. Micro Analysis
5.5. Long Generation Accuracy and Efficiency
We evaluate RetroInfer on the state-of-the-art reasoning model DeepSeek-R1-Distill-Llama-8B (guo2025deepseek, ) to test its performance on long-generation scenarios. We select several widely used benchmarks: AIME24 (aime24, ), AMC23 (amc23, ) and MATH-500 (lightman2023letsverifystepstep, ), which feature inputs with a few hundred tokens with long outputs ranging from 1K to 32K tokens. Following the standard evaluation settings (guo2025deepseek, ), we use $top\_p=0.95$, $top\_k=20$ and $temperature$ $=0.6$ during generation. Each test data is sampled eight times, and we report pass@8 as the accuracy metric. For the efficiency evaluation, we use input data from AIME and fix the output length to 16K tokens.
Since the input length is relatively short (less than 1K tokens), we do not construct the index during the prefilling stage. Instead, the prompt tokens are stored in the steady zone. As generation progresses, once the combined length of prompt and generated tokens exceeds 1K, we perform $k$-means clustering and initialize the index. Subsequent index updates follow the incremental update strategy described in §7. We set the update segment size to 1K to constrain the GPU memory usage and maintain the model accuracy.
As shown in Table 4, RetroInfer achieves comparable accuracy to full attention while delivering a 2.5$\times$ improvement in throughput, demonstrating the ability of RetroInfer on long-generation scenarios and the effectiveness of its index update strategy.
Table 4. Long Generation Accuracy (pass@8) and Efficiency (throughput@16K) on DeepSeek-R1-Distill-Llama-8B. RetroInfer matches the accuracy of full attention while delivering higher throughput.
Methods AIME24 AMC23 MATH-500 Throughput
Full attention 73.30 97.50 97.21 760.41
RetroInfer 76.70 95.00 96.41 1932.13